Water flows through a garden house with inner diameter 1.5 cm at a speed of 2.9 m/s. Calculate the speed of water escape from the nozzle if the opening is reduce to a diameter of 0.50 cm

1
Expert's answer
2022-01-25T06:49:55-0500

Solution

Using continuity equation of fluid mechanics

"A_1v_1=A_2v_2"

In form of diameter

"D_1^2v_1=D_2^2v_2"

Putting values diameters and speed

Velocity

"v_2=\frac{D_1^2 v_1}{D_2^2}\\=\frac{(0.015) ^2\times2.9}{(0.005) ^2}"

=26.1m/sec.